ADNAN MUHAMMED

Grade 12,

give examples of exothermic and endothermic reactions

give examples of exothermic and endothermic reactions

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear sachin,

Photosynthesis is an example of an endothermic chemical reaction. In this process, plants use the energy from the sun to convert carbon dioxide and water into glucose and oxygen. This reaction requires 15MJ of energy (sunlight) for every kilogram of glucose that is produced:

sunlight + 6CO2(g) + H2O(l) = C6H12O6(aq) + 6O2(g)

An example of an exothermic reaction is the mixture of sodium and chlorine to yield table salt. This reaction produces 411 kJ of energy for each mole of salt that is produced:

Na(s) + 0.5Cl2(s) = NaCl(s)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi